# GT-SVQ: A Linear-Time Graph Transformer for Node Classification Using Spiking Vector Quantization

### v1

@@ Abstract
Graph Transformers (GTs), which simultaneously integrate message passing and self-attention mechanisms, have achieved promising empirical results in some graph prediction tasks. Although these approaches show the potential of Transformers in capturing long-range graph topology information, issues concerning the quadratic complexity and high computing energy consumption severely limit the scalability of GTs on large-scale graphs. Recently, as brain-inspired neural networks, Spiking Neural Networks (SNNs), facilitate the development of graph representation learning methods with lower computational and storage overhead through the unique event-driven spiking neurons. Inspired by these characteristics, we propose a linear-time Graph Transformer using Spiking Vector Quantization (GT-SVQ) for node classification. GT-SVQ reconstructs codebooks based on rate coding outputs from spiking neurons, and injects the codebooks into self-attention blocks to aggregate global information in linear complexity. Spiking vector quantization effectively alleviates the issues concerning codebook collapse and the reliance on complex machinery (distance measure, auxiliary loss, etc.) in previous vector quantization-based graph learning methods. In experiments, we compare GT-SVQ with other state-of-the-art baselines on node classification datasets ranging from small to large. Experimental results show that GT-SVQ has achieved competitive performances on most datasets while maintaining up to 130× faster inference speed compared to other GTs. Our code is available at https://github.com/Zhhuizhe/GT-SVQ [^2] Work in progress.
1 Introduction
@@
Graph Transformers (GTs), as emerging graph representation learning paradigms, are proposed for alleviating inherent drawbacks of message passing neural networks like over-smoothing, over-squashing and local structure biases Oono and Suzuki (2019)Topping et al. (2021). Benefiting from the multi-head attention (MHA) modules, vanilla Transformers adaptively learn the global dependencies in input sequences without considering their distance Vaswani (2017). It also provides a solution for learning new topology among nodes while performing message aggregation on the graph data. Experiments demonstrate the immense potential of Transformers in handling global dependencies for graph data Rampášek et al. (2022)Bo et al. (2023). However, there is one critical drawback that Transformer with $O(N^{2})$ computation complexity is prohibitive for large-scale graphs. Furthermore, the all-pair similarity matrix leads to an increase in degrees of freedom, which often manifests as the full-size Transformers being highly prone to overfitting. Unlike observations in the field of computer vision or natural language process, previous studies show that eliminating redundant components and embracing a lightweight architecture like linear-time attention can significantly enhance the predictive performances of GTs Wu et al. (2022)Wu et al. (2024).
@@
Recently, with the development of neuromorphic computing, Spiking Neural Networks (SNNs) are poised to bridge the efficiency gap between elaborate network architectures and computation consumption. The defining feature of a SNN is the brain-like spiking mechanism which converts real-value signals into single-bit, sparse spiking signals based on its event-driven biological neurons. The single-bit nature enables us to adopt more addition operations rather than expensive multiply-and-accumulate operations on the spiking outputs, while the sparsity means spikes are cheap to be stored Eshraghian et al. (2023). These delightful characteristics have prompted some studies constructing lightweight graph representation learning frameworks to explore binary spike-based representations on the graph data Zhu et al. (2022)Li et al. (2023a)Li et al. (2023b). Previous experiments show that the potentiality of SNNs is still underestimated and underappreciated in the domain of graph representation learning. The role of SNNs in modeling the graph structure and dynamics warrants further investigation.
Figure 1: (left) Visualization results between full-precision and spike count node embeddings which are respectively generated by a message propagation method and spiking vector quantization on KarateClub. We highlight our observations in red. Notably, though some elements in full-precision embeddings are difficult to differentiate in color, their corresponding values are different. (right) Nodes are colored by their raw labels and spike count embeddings (also referred to as codewords) under the setting, $V_{th}=0.2$ and $T=4$. It shows that low-precision codewords implicitly group nodes based on learned class label information (e.g., nodes 13, 14, 16 and 24) and structural information (e.g., nodes 14, 15, 18 and 32). More implementation details and visualization results are attached in Appendix A.
We conduct preliminary experiments to uncover the role of spiking neurons in message propagation. As depicted in Figure1, there are three observations: (a) Those nodes within the same class tend to be closer in the embedding space by performing message propagation-based models. (b) The node embeddings from each propagation step are fed into spiking neurons to obtain spike count embeddings. It can be considered as generating the low-precision approximations of the original high-precision embeddings. (c) By converting the full-precision embeddings into spike count embeddings, nodes that are close in the original embedding space can be represented by the same low-precision embedding. It projects continuous high-precision representations into discrete low-precision representations, which sparks our curiosity about an interesting research question.
@@
Can we go beyond viewing spiking neurons as simple low-power units, and utilize discrete low-precision node representations to build more efficient Graph Transformers?
Deviating from previous works on efficient GTs, we consider spikes neurons as learnable quantizers, which links the rate coding outputs with the concept of vector quantization (VQ) Van Den Oord et al. (2017)Mentzer et al. (2023). We propose GT-SVQ, a linear-time Graph Transformer based on Spiking Vector Quantization for node classification. Specifically, GT-SVQ feeds full-precision embeddings from each message propagation into spiking neurons. Accumulating the spiking outputs to generate spike count embeddings enables nodes to be represented by a finite set of integer embeddings, which we also refer to as codewords. And the self-attention block guided by codewords aggregates the long-range information in linear complexity. It is worth noting that the above conversion from full-precision embeddings to codewords need not explicitly create a codebook. Different from prior works, considering spiking neurons as quantizers to generate codewords effectively addresses the codebook collapse defined as the under-usage of the codebook. The contributions of this paper are summarized as follows:
  • We conducted an in-depth analysis of spiking neurons in message propagation. Based on the observations, we propose Spiking Vector Quantization (SVQ), which provides a spike-driven learnable codebook paradigm to alleviate inherent issues in existing VQ-based graph learning methods.
  • We propose a linear-time Graph Transformer, GT-SVQ, based on the codebook guided self-attention mechanism. GT-SVQ takes the lead in introducing low-precision spike count embeddings from the narrower representation space into Graph Transformers. By calculating attention scores between nodes and node sets grouped by codewords, it effectively captures the long-range information with lower computational overhead.
@@
  • We conduct a comprehensive comparison with various state-of-the-art baselines across graphs with various scales. Extensive experiments show that GT-SVQ achieves competitive or even superior predictive performances on most datasets. Besides, GT-SVQ enjoys up to 130x faster inference speed compared to other GT baselines.
2 Preliminaries
Spiking Neural Network.
Although electrophysiological measurements can be accurately calculated by complex conductance-based neurons, the complexity limits their widespread deployment in deep neural networks. A simplified computational unit that retains biological characteristics, known as the Integrate-and-Fire neuron, has been proposed Salinas and Sejnowski (2002). IF neurons have three basic characteristics: Integrate, Fire and Reset. Firstly, the neuron integrates synaptic inputs from other neurons or external current $I$ to charge its cell membrane. Secondly, when the membrane potential reaches a pre-defined threshold value $V_{th}$, the neuron fires a spike $S$. Thirdly, the membrane potential of neuron will be reset to $V_{reset}$ after firing. The neuronal dynamics can be formulated as follows:
Integrate: $\;V^{t}=\Psi(V^{t-1},I^{t})=V^{t-1}+I^{t},$ (1)
Fire: $\;S^{t}=\Theta(V^{t}-V_{th})=\left\{\begin{array}[]{ll}1,&V^{t}-V_{th}\geq 0\\ 0,&otherwise\end{array},\right.$ (4)
Reset: $\;V^{t}=V^{t}(1-S^{t})+V_{reset}S^{t},$ (5)
where $V^{t}$ and $I^{t}$ denote the membrane potential and input current at time step $t$, respectively. $\Theta(\cdot)$ denotes the fire function, and the Heaviside step function is selected as the fire function in this paper. $\Psi(\cdot)$ is the membrane potential update function. Besides, there are two common variants of IF model, LIF and PLIF Gerstner et al. (2014)Fang et al. (2021b). The update function of these neurons can be formalized as follows:
LIF: $\;V^{t}=V^{t-1}+\frac{1}{\tau}(I^{t}-(V^{t-1}-V_{reset})),$ (6)
PLIF: $\;V^{t}=V^{t-1}+\frac{1}{1+exp(-a)}(I^{t}-(V^{t-1}-V_{reset})),$ (7)
where $\tau$ is the membrane time constant and $a$ is a trainable parameter, both of which are used to regulate how fast the membrane potential decays. Except for solely encoding input data into spikes, the rate coding scheme is widely adopted by SNNs to extract meaning from the spiking sequence. In this paper, we transform spiking outputs into spike counts that accumulate $S^{t}$ over $T$ time steps, $\hat{S}=\sum_{t=0}^{T}S^{t}$. Besides, we adopt surrogate gradients during error backpropagation to address the issue of zeros gradients caused by non-differentiable functions Neftci et al. (2019). The surrogate gradient method can be defined as $\Theta^{\prime}(x)\triangleq\theta^{\prime}(\alpha x)$, where $\alpha$ represents a smooth factor and $\theta(\cdot)$ represents a surrogate function Neftci et al. (2019).
@@ 2 Preliminaries
We represent a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is a set of nodes and $\mathcal{E}$ is a set of edges among these nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix of the graph. Let $N$ denotes the number of nodes. We define the $d$-dimension nodes’ attribute as $\mathbf{X}\in\mathbb{R}^{N\times d}$, which is known as the node feature matrix. For a given node $u\in\mathcal{V}$, GNN aggregates messages from its immediate neighborhood $N(u)$ and updates the node embedding $h_{u}$. This message passing process can be formulated as follows:
@@
where $h^{l}_{u}$ denotes the updated embedding of node $u$ in the $l$-th layer, $h^{l-1}_{u}$ is the embedding from the previous layer. $\operatorname{UPDATE}$ and $\operatorname{AGGREGATE}$ can be arbitrary differentiable functions.
@@
$Attn(\mathbf{X})=softmax(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d^{\prime}}})\mathbf{V},$ (9)
@@
where Query, Key and Value are calculated by learnable projection matrices $W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d^{\prime}}$. We omit the scalar factor for brevity.
3 GT-SVQ
@@
In this section, we comprehensively detail our approach, GT-SVQ. As depicted in Figure 2, GT-SVQ feeds the graph topology information into the Spiking Vector Quantization (SVQ) to map node embeddings to the spike count embeddings. These codewords will guide the aggregation process in the self-attention module. Besides, auxiliary message passing neural networks as position encoders provide node embeddings containing local positional information to the attention module. In what follows, we detail the implementation of SVQ (Section 3.1). Then, we highlight how the learnable codebook is introduced into self-attention to capture long-range information in the graph (Section 3.2). Finally, we revisit the entire architecture of GT-SVQ one by one (Section 3.3).
@@
Figure 2: The overview of GT-SVQ. Intuitively, the spiking vector quantization maps nodes to codewords which groups nodes by their structure and feature information. The codebook reconstructed from the spiking count embeddings is fed into the codebook guided self-attention. It assists the self-attention block in calculating attention scores from nodes to grouped node sets in linear time and memory complexity.
3.1 Spiking Vector Quantization
@@
Typically, spiking neural networks are designed for time-varying data. In order to meet the requirement for time-varying data, a common practice in spiking graph neural networks is to repeatedly pass the static graph to spiking neurons at each time step. It inevitably brings high computational and storage overheads. In this work, we collect the node embeddings over $T$ propagation steps as the sequential inputs required for spiking neurons, which effectively reduce additional overheads. Besides, we decompose spike count embeddings accumulated from spiking outputs to dynamically reconstruct codebooks. This avoids the issue of explicitly defining the codebook that contains a large number of unused codewords.
@@
Specifically, (i) we sample a $D$-dimension random feature matrix $\mathbf{R}\in\mathbb{R}^{N\times D}$ from a uniform distribution, and define a propagation operator $\mathbf{P}$. Based on $\mathbf{P}$ and $\mathbf{R}$, we update and collect node embeddings $\mathbf{M}\in\mathbb{R}^{N\times D}$ over $T$ propagation steps, =$\mathcal{M}=\{\mathbf{M}^{0},\mathbf{M}^{1},...,\mathbf{M}^{T}\}$. (ii) Given a SNN, it converts the above sequential node embeddings into spiking outputs $\mathcal{S}=\{\mathbf{S}^{0},\mathbf{S}^{1},...,\mathbf{S}^{T}\},\mathbf{S}\in\{0,1\}^{N\times D}$. (iii) The spike count embeddings $\hat{\mathbf{S}}\in\mathbb{Z}^{N\times D}$ are obtained by summing $\mathbf{S}^{t}$ over $T$ propagation steps. (iv) $\hat{\mathbf{S}}$ can be decomposed into the codebook $\mathbf{C}\in\mathbb{Z}^{B\times D}$ and the one-hot matrix $\mathbf{U}\in\{0,1\}^{N\times B}$. The above processes are defined as:
@@
where $Norm(\cdot)$ is the $l_{2}$ normalization, $\Theta(\cdot)$ and $\Phi(\cdot)$ is membrane potential update and fire functions. We choose similar propagation operators presented in previous work Eliasof et al. (2023). The codebook $\mathbf{C}$ is created by removing duplicate row vectors in $\hat{\mathbf{S}}$. The one-hot matrix $\mathbf{U}$ can be seen as the indices for where the codewords in $\hat{\mathbf{S}}$ ended up in $\mathbf{C}$. Hereafter, we will refer to $B$ as the size of the reconstructed codebook.
We define a discrete latent space $\tilde{\mathcal{C}}$ of the spiking count vector $\hat{s}\in\mathbb{Z}^{D}$ , $\forall\hat{s}\in\tilde{\mathcal{C}}$. The size of $\tilde{\mathcal{C}}$ is determined by both the number of propagation steps $T$ and the dimensionality of each random feature $D$, $|\tilde{\mathcal{C}}|=(T+1)^{D}$. This is because each dimension can take $T+1$ possible spike counts (considering the case where the total number of spikes is zero) across $T$ time steps. Different from vanilla VQ methods, which explicitly pre-define a codebook containing all latent embedding vectors for lookup, SVQ dynamically maps continuous full-precision embedding to the subset of $\tilde{\mathcal{C}}$ via the SNN. It enables the size of the reconstructed codebook to be much smaller, $B\ll|\tilde{\mathcal{C}}|$, while effectively alleviating the codebook collapse problem.
3.2 Codebook Guided Self-Attention
After the spiking vector quantization, we propose a codebook guided self-attention (CGSA) to capture long-range signals in linear complexity. Technically, node embeddings $\mathbf{H}\in\mathbb{R}^{N\times d^{\prime}}$ generated from auxiliary MPNNs are introduced as Query and Value. We materialize the Key $\hat{\mathbf{K}}$ based on the matrices $\mathbf{C}$ and $\mathbf{U}$. The attention functions are defined as:
@@ 3.2 Codebook Guided Self-Attention
where $d^{\prime}$ denotes the dimension of intermediate embeddings. In CGSA, $\hat{\mathbf{K}}$ is calculated based on the matrix multiplication between $\mathbf{G}$ and $\mathbf{U}$, rather than performing the nearest neighbor look-up on the entire codebook. It avoids the severe reliance on complex components like distance measures or auxiliary losses. Derived from Lingle (2023), the attention weights in eq 18 can be further factored:
@@
where $\mathbf{1}\in\mathbb{R}^{N}$. $\mathbf{U}^{T}\mathbf{1}=\{n_{b}\}^{B}_{b}\in\mathbb{Z}^{+}$ denotes the number of node embeddings in $\mathbf{C}$ mapped to the same codewords. The complexity of CGSA is $O(NBd_{v})$, where $B\ll N$. It can be considered that the computational overhead of CGSA grows linearly with the number of nodes. To avoid generating an excessively large codebook in the initial phase of learning, we perform a truncation strategy. We rank ${n_{b}}$ from high to low and select the top $B_{max}$ to generate a truncated codebook, which ensures the training efficiency of our model.
3.3 Overall Framework
As shown in Figure 2, the overview of GT-SVQ includes four modules: SVQ, auxiliary MPNN, CGSA and a classification head (CH). In the SVQ, we construct random features and spiking neurons for each layer. By defining a shared propagation operator, messages among nodes are collected and transformed into spike count embeddings $\hat{\mathbf{S}}$. Then auxiliary MPNNs as encoders generates node embeddings with local positional encodings. In the CGSA, the reconstructed codebook $\mathbf{C}$, codeword indices $\mathbf{U}$ and node embeddings $\mathbf{H}$ are fed into a linear-time self-attention. Different from the vanilla Transformer, we explicitly inject graph inductive bias by coding global structural information into spike count embeddings. These four parts can be written as follows:
$\mathbf{U}^{l},\mathbf{C}^{l}$ $=\operatorname{SVQ}^{l}(\mathbf{A}),$ (20)

### v2

# GT-SNT: A Linear-Time Transformer for Large-Scale Graphs via Spiking Node Tokenization

@@ Abstract
Graph Transformers (GTs), which integrate message passing and self-attention mechanisms simultaneously, have achieved promising empirical results in graph prediction tasks. However, the design of scalable and topology-aware node tokenization has lagged behind other modalities. This gap becomes critical as the quadratic complexity of full attention renders them impractical on large-scale graphs. Recently, Spiking Neural Networks (SNNs), as brain-inspired models, provided an energy-saving scheme to convert input intensity into discrete spike-based representations through event-driven spiking neurons. Inspired by these characteristics, we propose a linear-time Graph Transformer with Spiking Node Tokenization (GT-SNT) for node classification. By integrating multi-step feature propagation with SNNs, spiking node tokenization generates compact, locality-aware spike count embeddings as node tokens to avoid predefined codebooks and their utilization issues. The codebook guided self-attention leverages these tokens to perform node-to-token attention for linear-time global context aggregation. In experiments, we compare GT-SNT with other state-of-the-art baselines on node classification datasets ranging from small to large. Experimental results show that GT-SNT achieves comparable performances on most datasets and reaches up to 130× faster inference speed compared to other GTs.
Codehttps://github.com/Zhhuizhe/GT-SNT
Introduction
@@
Graph Transformers (GTs), as emerging graph representation learning paradigms, are proposed for alleviating inherent drawbacks of message passing neural networks like over-smoothing, over-squashing and local structure biases (oono2019graph; topping2021understanding). Benefiting from the multi-head attention (MHA) modules, vanilla Transformers adaptively learn the global dependencies of input sequences (vaswani2017attention). It also provides a solution for learning new topology among nodes while performing message aggregation on the graph data. Experiments demonstrate the immense potential of Transformers in handling global or long-range interactions (rampavsek2022recipe; bo2023specformer). However, there is one critical drawback that Transformers with $O(N^{2})$ computation complexity are prohibitive for large-scale graphs. Although empirical experiments show that designing linear-time attention in GTs can significantly reduce redundant computational cost (wu2022nodeformer; wu2024sgformer), such approaches introduce potential issues related to over-globalization (xing2024less). As an orthogonal technique, encapsulating structural or semantic information into node tokens provides a promising direction that enables GTs to efficiently capture salient graph properties via complementary local and global tokens (wang2025gqt; luo2024nodeid).
@@
Recently, with the development of neuromorphic computing, Spiking Neural Networks (SNNs) offer a promising path toward energy-saving neural networks that achieve competitive performances using as few computations as possible. The defining feature of SNNs is the brain-like spiking mechanism which converts real-value signals into single-bit, sparse spiking signals based on event-driven biological neurons. The single-bit nature enables us to adopt more addition operations rather than expensive multiply-and-accumulate operations on the spiking outputs. Besides, the sparsity means spikes are cheap to store (eshraghian2023training). These delightful characteristics have prompted some studies constructing lightweight Spiking Graph Neural Networks (SGNNs) to explore spike-based representations on the graph data (zhu2022spiking; li2023scaling; li2023graph). Despite tremendous progress in replacing artificial neurons with spiking counterparts, the broader use of SNNs for graph representation learning remains underexplored. Existing SGNNs tend to focus on their low-power advantages, while overlooking the expressive power of their sparse and discrete embeddings. SNNs naturally project continuous, high-precision inputs into low-precision, event-driven representations. This observation sparks our curiosity about an interesting research question:
@@
Can we go beyond viewing spiking neurons merely as simple low-power units, and utilize spiking representations to build efficient tokenized Graph Transformers?
Figure 1: Frameworks of different tokenization strategies. Spiking Node Tokenization converts sequential inputs into spike count embeddings via spiking neurons, which yields a codebook-free, locality-aware node tokenization.
Tokenizers enable Transformer-based architectures to efficiently learn high-level representations from a small and manageable number of tokens. As shown in Figure 1, an exemplary implementation of tokenizers is VQ-VAE (van2017neural), which retrieves codewords from a fixed codebook to replace original embeddings. To alleviate the issue of under-utilization, Lookup Free Quantization (LFQ) (yu2023language) proposes a learnable scalar quantization mechanism to build an implicit codebook. Inspired by existing tokenization strategies, we argue that SNNs inherently convert sequential inputs into sparse and binary spike trains, which not only reduces energy consumption but also offers a novel sequence-to-token approach. An energy-efficient node tokenization driven by SNNs is demonstrated in Figure 1. Spiking Node Tokenization (SNT) feeds full-precision embeddings from each propagation into spiking neurons. Spike count representations accumulated from spiking outputs represent nodes by a finite set of integer vectors, which we refer to as codewords.
Built upon the spike-driven node tokenizer, we propose GT-SNT, a linear-time Graph Transformer based on Spiking Node Tokenization for large-scale graphs. Specifically, GT-SNT balances local and global representations via two key components: Spiking Node Tokenization (SNT) and Codebook Guided Self-Attention (CGSA). SNT converts random feature propagation sequences into discrete, locality-aware node tokens. CGSA computes node-to-token attention scores based on generated spike count representations to capture long-range interactions in linear complexity. GT-SNT features several advantages: a) Different from existing tokenized GTs, SNT proposes a generative and adaptive tokenization scheme built upon spiking neurons to evade the issue of codebook collapse. b) SNT encodes graph inductive bias through well-designed sequential data. It leverages rich spiking dynamics and enables locality-aware quantization beyond simple binary quantization. c) CGSA is an efficient attention module by dynamically deriving the Key matrix from spike count representations without introducing complex machinery (e.g., distance metrics or auxiliary losses). The contributions of this paper are summarized as follows:
  • We propose Spiking Node Tokenization, which links SNNs with the sequential positional encoding data to generate locality-aware node tokens. To the best of our knowledge, our work is the first to explore the usage of SNN as an energy-saving tokenizer for graph data.
  • We tailor a node-to-token attention module with linear complexity, CGSA, which captures global topological information upon spike count node tokens. It provides a straightforward global aggregation guided by node tokens to balance both efficiency and expressiveness.
@@
  • We conduct a comprehensive comparison with state-of-the-art baselines across graphs with various scales. Extensive experiments show that GT-SNT achieves comparable or even superior predictive performance on most datasets. Besides, GT-SNT enjoys up to 130x faster inference speed compared to other GT baselines.
Preliminaries
@@
We represent a graph as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is a set of nodes and $\mathcal{E}$ is a set of edges among these nodes, $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the adjacency matrix of the graph. Let $N$ denote the number of nodes. We define the $d$-dimension nodes’ attribute as $\mathbf{X}\in\mathbb{R}^{N\times d}$, which is known as the node feature matrix. For a given node $u\in\mathcal{V}$, GNN aggregates messages from its immediate neighborhood $N(u)$ and updates the node embedding $h_{u}$. This message passing process can be formulated as follows:
@@
where $h^{l}_{u}$ denotes the updated embedding of node $u$ in the $l$-th layer. $h^{l-1}_{u}$ is the embedding from the previous layer. $\operatorname{UPDATE}$ and $\operatorname{AGGREGATE}$ can be arbitrary differentiable functions.
@@
$\operatorname{Attn}(\mathbf{X})=\operatorname{softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d^{\prime}}})\mathbf{V},$ (2)
@@
where Query, Key and Value are calculated by learnable projection matrices $W_{q},W_{k},W_{v}\in\mathbb{R}^{d\times d^{\prime}}$. We omit the scaling factor hereafter for brevity.
Spiking Neural Network.
Although electrophysiological measurements can be accurately calculated by complex conductance-based neurons, the complexity limits their widespread deployment in deep neural networks. A simplified computational unit, which is known as the Integrate-and-Fire neuron, has been proposed (salinas2002integrate). IF neurons have three basic characteristics: Integrate, Fire and Reset. Firstly, the neuron integrates synaptic inputs from other neurons or external current $I$ to charge its cell membrane. Secondly, when the membrane potential reaches a pre-defined threshold value $V_{th}$, the neuron fires a spike $S$. Thirdly, the membrane potential of the neuron will be reset to $V_{reset}$ after firing. The neuronal dynamics can be formulated as follows:
Integrate: $\;V^{t}=\Psi(V^{t-1},I^{t})=V^{t-1}+I^{t},$ (4)
Fire: $\;S^{t}=\Theta(V^{t}-V_{th})=\left\{\begin{array}[]{ll}1,&V^{t}-V_{th}\geq 0\\ 0,&\text{otherwise}\end{array},\right$. (7)
Reset: $\;V^{t}=V^{t}(1-S^{t})+V_{reset}S^{t},$ (8)
where $V^{t}$ and $I^{t}$ denote the membrane potential and input current at time step $t$, respectively. $\Theta(\cdot)$ denotes the fire function, implemented as a Heaviside step function. $\Psi(\cdot)$ is the membrane potential update function. Besides, there are two common IF variants, LIF and PLIF (gerstner2014neuronal; fang2021incorporating). The update function of these neurons can be formalized as follows:
LIF: $\;V^{t}=V^{t-1}+\frac{1}{\tau}(I^{t}-(V^{t-1}-V_{reset})),$ (9)
PLIF: $\;V^{t}=V^{t-1}+\frac{1}{1+\exp(-\beta)}(I^{t}-(V^{t-1}-V_{reset})),$ (10)
where $\tau$ is the membrane time constant and $\beta$ is a trainable parameter. They are used to regulate how fast the membrane potential decays. The rate coding scheme is widely adopted by SNNs to extract spiking dynamics from sequential data. In this paper, we transform spiking outputs into spike counts that accumulate $S^{t}$ over $T$ time steps, $\hat{S}=\sum_{t=0}^{T}S^{t}$. Besides, we adopt surrogate gradients during error backpropagation to address the issue of zero gradients caused by non-differentiable functions (neftci2019surrogate). The surrogate gradient method can be defined as $\Theta^{\prime}(x)\triangleq\theta^{\prime}(\alpha x)$, where $\alpha$ represents a smooth factor and $\theta(\cdot)$ represents a surrogate function (neftci2019surrogate).
GT-SNT
@@
In this section, we comprehensively detail our approach, GT-SNT. As depicted in Figure 2, GT-SNT feeds the intermediate embeddings generated from feature propagation into the Spiking Node Tokenization (SNT) to inject graph positional information into the spike count embeddings. These codewords will guide the aggregation process in the self-attention module. Additionally, auxiliary message passing neural networks serve as encoders that encapsulate both node features and local graph topology to the attention module. In what follows, we detail the implementation of SNT. Then, we highlight how CGSA dynamically reconstructs a codebook only containing used codewords to apply node-to-token attention on large-scale graphs. Finally, we revisit the entire architecture of GT-SNT one by one.
@@
Figure 2: The overview of GT-SNT. Intuitively, the spiking node tokenization maps nodes to spike count vectors, which essentially groups nodes by their tokens. The codebook reconstructed from the spike count embeddings is fed into the codebook guided self-attention. It assists the self-attention block in calculating attention scores from nodes to grouped node sets in linear time and memory complexity.
Spiking Node Tokenization
@@
Typically, spiking neural networks are designed for time-varying data. In order to meet the requirement for time-varying data, a common practice in spiking graph neural networks is to repeatedly pass the static graph to spiking neurons at each time step. It inevitably brings high computational and storage overheads. In this work, we collect the node embeddings over $T$ propagation steps as the sequential inputs required for spiking neurons, which effectively reduce additional overheads. Besides, we decompose spike count embeddings accumulated from spiking outputs to dynamically reconstruct codebooks. This avoids the issue of explicitly defining the entire codebook that contains a large number of unused codewords.
@@
Specifically, (i) we sample a $D$-dimension learnable random feature matrix $\mathbf{R}\in\mathbb{R}^{N\times D}$ from a uniform distribution, and define a propagation operator $\mathbf{P}$. Based on $\mathbf{P}$ and $\mathbf{R}$, we update and collect node embeddings $\mathbf{M}\in\mathbb{R}^{N\times D}$ over $T$ propagation steps, $\mathcal{M}=\{\mathbf{M}^{0},\mathbf{M}^{1},...,\mathbf{M}^{T}\}$. (ii) Given a SNN, it converts the above sequential node embeddings into spiking outputs $\mathcal{S}=\{\mathbf{S}^{0},\mathbf{S}^{1},...,\mathbf{S}^{T}\},\mathbf{S}\in\{0,1\}^{N\times D}$. (iii) The spike count embeddings $\hat{\mathbf{S}}\in\mathbb{Z}^{N\times D}$ are obtained by summing $\mathbf{S}^{t}$ over $T$ propagation steps. (iv) $\hat{\mathbf{S}}$ can be decomposed into the codebook $\mathbf{C}\in\mathbb{Z}^{B\times D}$ and the one-hot matrix $\mathbf{U}\in\{0,1\}^{N\times B}$. The above processes are defined as:
@@
where $Norm(\cdot)$ normalizes output messages to the range of threshold membrane potential. $\Theta(\cdot)$ and $\Phi(\cdot)$ are membrane potential update and fire functions, respectively. We choose similar propagation operators presented in previous work (eliasof2023graph). The codebook $\mathbf{C}$ is created by removing duplicate row vectors in $\hat{\mathbf{S}}$. The one-hot matrix $\mathbf{U}$ can be seen as the indices for where the codewords in $\hat{\mathbf{S}}$ ended up in $\mathbf{C}$. Hereafter, we will refer to $B$ as the size of the reconstructed codebook.
[…]
Codebook Guided Self-Attention
After the Spiking Node Tokenization, we propose a codebook guided self-attention (CGSA) to capture long-range signals in linear complexity. Specifically, node embeddings $\mathbf{H}\in\mathbb{R}^{N\times d^{\prime}}$ from auxiliary MPNNs are introduced as Query and Value. We materialize the Key $\hat{\mathbf{K}}$ based on the matrices $\mathbf{C}$ and $\mathbf{U}$. The attention functions are defined as:
@@
where $d^{\prime}$ denotes the dimension of intermediate embeddings. In CGSA, $\hat{\mathbf{K}}$ is calculated based on the matrix multiplication between $\mathbf{G}$ and $\mathbf{U}$, rather than performing the nearest neighbor look-up on the entire codebook. It avoids the severe reliance on complex components like distance measures or auxiliary losses. Derived from (lingle2023transformer), the attention weights in Eq 18 can be further factored:
@@
where $\mathbf{1}\in\mathbb{R}^{N}$. $\mathbf{U}^{T}\mathbf{1}=\{n_{b}\}^{B}_{b}\in\mathbb{Z}^{+}$ denotes the number of spike count embeddings in $\hat{\mathbf{S}}$ mapped to the same codewords. The complexity of CGSA is $O(NBd_{v})$, where $B\ll N$. It can be considered that the computational overhead of CGSA grows linearly with the number of nodes. To avoid generating an excessively large codebook in the initial phase of learning, we perform a truncation strategy. We rank $n_{b}$} from high to low and select the top $B_{max}$ to generate a truncated codebook, which ensures the training efficiency of GT-SNT.
Overall Framework
As shown in Figure 2, the overview of GT-SNT includes four modules: SNT, auxiliary MPNN, CGSA and a classification head (CH). In the SNT, we construct learnable random features and spiking neurons for each layer. By defining a shared propagation operator, messages among nodes are collected and transformed into spike count embeddings $\hat{\mathbf{S}}$. Then auxiliary MPNNs as encoders generate node embeddings with semantic and local neighborhood information. In the CGSA, the reconstructed codebook $\mathbf{C}$, codeword indices $\mathbf{U}$ and node embeddings $\mathbf{H}$ are fed into a linear-time self-attention to capture global topological information. Different from the vanilla Transformer, CGSA employs global node-to-token attention to actively introduce graph inductive bias. These four parts can be written as follows:
$\mathbf{U}^{l},\mathbf{C}^{l}$ $=\operatorname{SNT}^{l}(\mathbf{A}),$ (20)